# The super learner for time-to-event outcomes: A tutorial

### v1

@@ 6.3 Methods: application and evaluation in the test data
To evaluate the performance of the SLs in the test dataset we use a range of measures of discrimination, calibration, and overall performance, guided by the recommendations of McLernon and others (2022). We estimate the Brier score at time 10 years and the scaled Brier score, which compares the estimated Brier score from the model to the Brier score obtained if everyone is assigned the same survival probability obtained from a Kaplan-Meier analysis (Graf and others, 1999). To assess discrimination we estimate the C-index up to time 10, and the cumulative dynamic AUCt at time 10. For the C-index we use Uno’s C (Uno and others, 2011) which accommodates censoring through weights. Our code also shows how to obtain calibration plots, but we do not show these here.
All of the measures used to evaluate the predictive performance require inverse probability for censoring weights, which are estimated using a model for the censoring distribution fitted in the test data. For this it is commonly assumed that censoring is independent, with censoring weights being estimated as the inverse of the Kaplan-Meier estimates for the censoring survival probabilities. We apply this approach. In the example code we also illustrate estimation of the censoring weights using a SL for the censoring distribution, fitted in the test data.
@@ 6.4 Results
Table 1 shows the estimated measures of predictive performance using the individual (continuous-time) learners, the discrete-time ensemble SL, the continuous-time ensemble SL, and the continuous-time non-ensemble SL. Tables S1, S2, S3, S4 show the coefficients (or weights) attached to different candidate learners in the discrete-time and continuous-time ensemble SLs, based on SLs using simple and richer sets of candidate learners.
@@
Of the individual learners the random survival forest had the best predictive performance, in terms of having the lowest Brier sore (and highest scaled Brier score) and the best discrimination. The standard Cox regression had the worst predictive performance overall.
@@
In the discrete-time SL, the version in which we discretised using 100 periods had worse overall performance, as measured by the Brier score, compared to the version in which we discretised using only 10 time periods. However the discrimination measures were very similar using the two levels of time discretisation. The difference is likely to be in part due to poor modelling of the baseline hazard in the version using 100 time periods. Recall that when using 10 time periods we included an indicator for each time period in each candidate learner, whereas when using 100 time periods we included linear, quadratic and cubic terms for time. The discrete-time SL using 100 time periods had worse overall performance than using the Cox model as an individual learner, albeit better discrimination.
@@
The continuous-time SLs gave better performance than the discrete-time SLs, in particular when looking at overall performance measured using the Brier score. The continuous-time non-ensemble SL of Munch and Gerds (2024) selected the random survival forest as the best model for the event time distribution, and hence the results are the same as shown when using the random survival forest as an individual learner. The continuous-time ensemble SL of Westling and others (2023) gave almost identical predictive performance to the continuous-time non-ensemble SL (and the random survival forest).
@@
Using a richer set of candidate learners gave better predictive performance, though the gains were modest in this example (Table 1). When implemented using a simple set of candidate learners, the discrete-time SL with 10 time periods placed most weight on the GLM with Lasso and on the random forest using both the L2 and loglik loss functions (Table S1). The GLM with Lasso was the single best performing model using both loss functions. The version using the L2 loss function placed zero weight on the standard GLM, whereas the version using the loglik loss function gave that model a small weight. By contrast, the discrete-time SL with 100 time periods and using a simple set of candidate learners placed most weight on the GLM and zero weight on the GLM with Lasso, for both loss functions. With 100 time periods and using the L2 loss function, the single best performing model was the GLM with Lasso, however the standard GLM has very similar performance, which can explain why the GLM with Lasso was assigned zero weight in the ensemble. With 100 time periods and using the loglik loss function, the single best performing model was the standard GLM. When we used a richer set of candidate learners (Table S3), the discrete-time SL with 10 time periods placed most weight on a GAM with 3 degrees of freedom, but also placed non-zero weight on the mean estimator, the GLM with Lasso, a GAM with 2 degrees of freedom, one XGBoost learner (out of 7), and one random forest (out of 3).
@@
The continuous-time ensemble SL using the simpler set of candidate learners assigned most weight to the random survival forest for the event distribution (Table S2), some weight to the Cox model with Lasso, and zero weight to the Kaplan-Meier or the standard Cox model. When using a richer set of learners (Table S4), all of the weight was shared between the Cox regression with continuous covariates modelled using GAMs and the random survival forest. Corresponding results for the censoring distribution are shown in Tables S2 and S4.
@@
Implementations of the discrete-time approach and the continuous-time approach of Munch and Gerds (2024) performed ‘by hand’ using a simple set of candidate learners gave similar results as found using the packages (see Supplementary Table S5).
@@
Super learners: simple set of learners (Tables S1, S2)
@@
Discrete-time SL: 10 periods, loglik loss 0.203 17.6 71.9 74.4
@@
Discrete-time SL: 100 periods, loglik loss 0.213 13.5 71.6 74.8
@@
Continuous-time SL: Munch & Gerds 2024 0.196 20.6 71.7 75.8
Super learners: richer set of learners (Tables S3, S4)
@@
Discrete-time SL: 10 periods, loglik loss 0.198 19.5 72.3 75.2
@@ 7 Discussion
The SL is a powerful method for enabling flexible estimation of conditional expected outcome functions, which selects a best single model from a set of candidate learners or derives an optimal combination or ‘ensemble’. The candidate learners can include parametric and machine learning models, and the ensemble SL performs asymptotically at least as well as the single best performing model (van der Laan and others, 2007; Polley and others, 2011). In many medical contexts the interest lies in a time-to-event outcome, and specifically in estimating or predicting a survival probability or a risk conditional on a set of individual characteristics, with estimation challenged by censoring. SLs for time-to-event outcomes have been developed, with new implementations appearing in recent years. However, the literature on the SL for time-to-event outcomes is technical, and somewhat disjointed, and a reader may find it challenging to gather together the full details of how these methods work and can be implemented based on the current literature. In this paper we have provided a general overview of the SL for time-to-event outcomes, followed by details of three specific implementations covering discrete-time and continuous-time versions of the SL. We have given a summary of how the methods can be implemented in R and illustrated the methods using an open access data set. R code is provided at https://github.com/ruthkeogh/superlearner_survival_tutorial, enabling others to reproduce our results. This includes by-hand code implementations that may help to provide additional clarification as well as helping users to adapt the methods in ways that may not be covered by the existing packages.
@@
Munch and Gerds (2024) noted that a drawback of their approach is that the loss function targets the conditional state occupation probabilities rather than the conditional survival probability, which is the target estimand. However their simulation studies indicated that good performance for the former translates into good performance for the latter. The loss functions used by Westling and others (2023) directly target the conditional survival probability, however Munch and Gerds (2024) pointed out that no general theoretical guarantees appear to have been given for the procedure of Westling and others (2023). Munch and Gerds (2024) compared their approach with that of Westling and others (2023) in a simulation study, finding that the two methods performed similarly, with their state learner approach often having slightly better performance. An important advantage of the state learner approach of Munch and Gerds (2024) is that it extends to handle competing events, with full details of this provided in the original paper.
@@
An alternative continuous-time SL was proposed by Golmakani and Polley (2020). However, their approach is restricted to candidate learners that involve a linear predictor and have a proportional hazards form. It therefore excludes methods such as random survival forests. Their ensemble step is also somewhat more complicated than those explained for the methods summarised in this tutorial. For these reasons we decided to exclude this approach from our overview. Their methods can be implemented using code available at https://github.com/kgolmakani/SuperLearner-Survival.git and https://github.com/kgolmakani/survSL.git. There is another R package implementing a SL called survivalSL (Foucher and Sabathe, 2025), however we did not find detailed documentation at the time of writing. In terms of implementation, the SuperLearner package (Polley and others, 2024) is currently the most well developed and flexible. The survsuperlearner package and the statelearner code of Munch and Gerds (2024) currently have some restrictions in terms of the candidate learners they can accommodate.
@@
In this paper we have focused on SL methods that target estimation of a discrete-time hazard, a survival function or a state occupation probability. An alternative approach is to target pseudo observations (Andersen and others, 2003), which enable censored survival data to be analysed as continuous outcomes. Sachs and others (2019) described a SL using pseudo observations, which enables use of models for continuous outcomes in the set of candidate learners. They focused on a setting with competing risks and proposed a loss function based on an AUC. After calculating the pseudo observations they used the SuperLearner package with the pseudo observations as a continuous outcomes, using a bespoke loss function. Details of the implementation are provided at https://github.com/sachsmc/pseupersims. An application of this approach to our Rotterdam data example is summarised in Supplementary Section S1. Cwiling and others (2024) also considered a SL based on pseudo observations, focusing on estimation of restricted mean survival times using a quadratic loss function. A drawback of using pseudo observations in prediction is that it can give risk estimates that are negative or greater than 1. The prediction model also has to be estimated separately for different time horizons $\tau$, which can lead to survival probabilities that not monotonically decreasing over time.
@@
In our implementations of the methods our aim was not to conduct a formal comparison, but instead to illustrate how they can be applied. We considered using both a relatively simple set of candidate learners and a richer set of learners. In practice, using a diverse set of learners is a good idea, as it results in lower correlation between the errors from different models, which is desirable (Brown and others, 2005). For example, Wey and others (2015) suggested that including lots of similar models (e.g different Cox models, with different ways of entering the covariates) is not advisable, as the predictions from these models will be highly correlated. An overview of machine learning algorithms for survival data is provided by Sonabend and Bender (2025). Some simulation based comparisons of the methods discussed in this paper have been conducted elsewhere (Munch and Gerds, 2024; Wolock and others, 2024). In practice, the choice between different methods, and of the candidate learners to be included, is likely to be based on practical considerations, including computational.
@@
Table S1: Coefficients (or weights) for the candidate learners (based on a simple set of candidate learners) for discrete-time SL, for 10 or 100 time periods and using the L2 or loglik loss function.
Candidate learner 10 periods, L2 loss 10 periods, loglik loss 100 periods, L2 loss 100 periods, loglik loss
@@
Table S3: Coefficients (or weights) for the candidate learners (based on a richer set of candidate learners) for discrete-time SL, for 10 time periods and using the L2 loss function.
@@
Discrete-time SL: 100 periods, L2 loss [by hand] 0.213 13.5 72.1 75.3
Discrete-time SL: 100 periods, loglik loss [by hand] 0.201 18.3 71.9 74.8
Discrete-time SL: 100 periods, ipcw loss [by hand] 0.224 9.1 70.9 73.7
S1 SL based on pseudo observations (Sachs and others, 2019)

### v2

@@ 6.3 Methods: application and evaluation in the test data
To evaluate the performance of the SLs in the test dataset we use a range of measures of discrimination, calibration, and overall performance, guided by the recommendations of McLernon and others (2022). We estimate the Brier score at time 10 years and the scaled Brier score, which compares the estimated Brier score from the model to the Brier score obtained if everyone is assigned the same survival probability obtained from a Kaplan-Meier analysis (Graf and others, 1999). To assess discrimination we estimate the C-index up to time 10 years, and the cumulative dynamic AUCt at time 10 years. For the C-index we use Uno’s C (Uno and others, 2011) which accommodates censoring through weights. Our code also shows how to obtain calibration plots, but we do not show these here.
All of the measures used to evaluate the predictive performance require inverse probability of censoring weights, which are estimated using a model for the censoring distribution fitted in the test data. For this it is commonly assumed that censoring is independent, with censoring weights being estimated as the inverse of the Kaplan-Meier estimates for the censoring survival probabilities. We apply this approach. In the example code we also illustrate estimation of the censoring weights using a SL for the censoring distribution, fitted in the test data.
@@ 6.4 Results
Table 1 shows the estimated measures of predictive performance using the individual (continuous-time) learners, the discrete-time ensemble SL, the continuous-time ensemble SL, and the continuous-time non-ensemble SL. Supplementary Tables S1-S4 show the coefficients (or weights) attached to different candidate learners in the discrete-time and continuous-time ensemble SLs, based on SLs using simple and richer sets of candidate learners.
@@
Of the individual learners the random survival forest gives the best predictive performance, in terms of having the lowest Brier score, highest scaled Brier score, and the best discrimination. The standard Cox regression has the worst predictive performance overall.
@@
In the discrete-time SL, the version in which we discretised using 100 periods has worse overall performance, as measured by the Brier score, compared to the version in which we discretised using only 10 time periods. However the discrimination measures are very similar using the two levels of time discretisation. The difference is likely to be in part due to poor modelling of the baseline hazard in the version using 100 time periods. Recall that when using 10 time periods we include an indicator for each time period in each candidate learner, whereas when using 100 time periods we include linear, quadratic and cubic terms for time. The discrete-time SL using 100 time periods has worse overall performance than using the Cox model as an individual learner, but better discrimination.
@@
The continuous-time SLs give better performance than the discrete-time SLs, in particular when looking at overall performance measured using the Brier score. The continuous-time non-ensemble SL of Munch and Gerds (2025) selects the random survival forest as the best model for the event time distribution, and hence the results are the same as shown when using the random survival forest as an individual learner. The continuous-time ensemble SL of Westling and others (2023) gives almost identical predictive performance to the continuous-time non-ensemble SL (and the random survival forest).
@@
Using a richer set of candidate learners gives better predictive performance, though the gains are modest in this example (Table 1). When implemented using a simple set of candidate learners, the discrete-time SL with 10 time periods places most weight on the GLM with Lasso and on the random forest using both the $L2$ and loglikloglik loss functions (Supplementary Table S1). The GLM with Lasso was the single best performing model using both loss functions. The version using the $L2$ loss function places zero weight on the standard GLM, whereas the version using the loglikloglik loss function gives that model a small weight. By contrast, the discrete-time SL with 100 time periods and using a simple set of candidate learners places most weight on the GLM and zero weight on the GLM with Lasso, for both loss functions. With 100 time periods and using the $L2$ loss function, the single best performing model is the GLM with Lasso, however the standard GLM has very similar performance, which can explain why the GLM with Lasso was assigned zero weight in the ensemble. With 100 time periods and using the loglikloglik loss function, the single best performing model was the standard GLM. When we use a richer set of candidate learners (Supplementary Table S3), the discrete-time SL with 10 time periods places most weight on a GAM with 3 degrees of freedom, but also places non-zero weight on the mean estimator, the GLM with Lasso, a GAM with 2 degrees of freedom, one XGBoost learner (out of 7), and one random forest (out of 3).
@@
The continuous-time ensemble SL of Westling and others (2023) using the simpler set of candidate learners assigns the most weight to the random survival forest for the event distribution (Supplementary Table S2), some weight to the Cox model with Lasso, and zero weight to the Kaplan-Meier or the standard Cox model. When using a richer set of learners (Supplementary Table S4), all of the weight is shared between the Cox regression with continuous covariates modelled using GAMs and the random survival forest. Corresponding results for the censoring distribution are shown in Supplementary Tables S2 and S4.
@@
Implementations of the discrete-time approach and the continuous-time approach of Munch and Gerds (2025) performed ‘by hand’ using a simple set of candidate learners give similar results as found using the packages (see Supplementary Table S5).
@@
Super learners: simple set of learners (see Supplementary Tables S1 and S2)
@@
Discrete-time SL: 10 periods, loglikloglik loss 0.203 17.6 71.9 74.4
@@
Discrete-time SL: 100 periods, loglikloglik loss 0.213 13.5 71.6 74.8
@@
Continuous-time SL: Munch & Gerds 2025 0.196 20.6 71.7 75.8
Super learners: richer set of learners (see Supplementary Tables S3 and S4)
@@
Discrete-time SL: 10 periods, loglikloglik loss 0.198 19.5 72.3 75.2
@@ 7 Discussion
The SL is a powerful method for enabling flexible estimation of conditional expected outcome functions, which selects a best single model from a set of candidate learners or derives an optimal combination or ‘ensemble’. The candidate learners can include parametric and machine learning models, and the ensemble SL performs asymptotically at least as well as the single best performing model (van der Laan and others, 2007; Polley and others, 2011). In many medical contexts the interest lies in a time-to-event outcome, and specifically in estimating or predicting a survival probability or a risk conditional on a set of individual characteristics, with estimation challenged by censoring. SLs for time-to-event outcomes have been developed, with new implementations appearing in recent years. However, the literature on the SL for time-to-event outcomes is technical, and somewhat disjointed, and a reader may find it challenging to gather together the full details of how these methods work and can be implemented based on the current literature. In this tutorial paper we have provided a general overview of the SL for time-to-event outcomes, followed by details of three specific implementations covering discrete-time and continuous-time versions of the SL. We have given a summary of how the methods can be implemented in R and illustrated the methods using an open access data set. R code is provided at https://github.com/ruthkeogh/superlearner_survival_tutorial, enabling others to reproduce our results. This includes by-hand code implementations that may help to provide additional clarification as well as helping users to adapt the methods in ways that may not be covered by the existing packages.
@@
Munch and Gerds (2025) noted that a drawback of their joint survival SL approach is that the loss function targets the conditional state occupation probabilities rather than the conditional survival probability, which is the target estimand. However their simulation studies indicated that good performance for the former translates into good performance for the latter. The loss functions used by Westling and others (2023) directly target the conditional survival probability, however Munch and Gerds (2025) pointed out that no general theoretical guarantees appear to have been given for the procedure of Westling and others (2023). Munch and Gerds (2025) compared their approach with that of Westling and others (2023) in a simulation study, finding that the two methods performed similarly, with their joint survival SL approach often having slightly better performance. An important advantage of the approach of Munch and Gerds (2025) is that it extends to handle competing events, with details of this provided in their paper.
@@
An alternative continuous-time SL was proposed by Golmakani and Polley (2020). However, their approach is restricted to candidate learners that involve a linear predictor and have a proportional hazards form. It therefore excludes methods such as random survival forests. Their ensemble step is also somewhat more complicated than those explained for the methods summarised in this tutorial. For these reasons we decided to exclude this approach from our overview. Their methods can be implemented using code available at https://github.com/kgolmakani/SuperLearner-Survival.git and https://github.com/kgolmakani/survSL.git. There is another R package implementing a SL called survivalSL (Foucher and Sabathe, 2025), however we did not find detailed documentation at the time of writing. In terms of implementation, the SuperLearner package (Polley and others, 2024) is currently the most well developed and flexible.
@@
In this paper we have focused on SL methods that are based on estimation of a discrete-time hazard, a survival function or a state occupation probability. An alternative approach is to target pseudo observations (Andersen and others, 2003), which enable censored survival data to be analysed as continuous outcomes. Sachs and others (2019) described a SL using pseudo observations, which enables use of models for continuous outcomes in the set of candidate learners. They focused on a setting with competing risks and proposed a loss function based on an AUC. After calculating the pseudo observations they used the SuperLearner package with the pseudo observations as a continuous outcomes, using a bespoke loss function. Details of the implementation are provided at https://github.com/sachsmc/pseupersims. An application of this approach to our Rotterdam data example is summarised in Supplementary Section S1. Cwiling and others (2025) also considered a SL based on pseudo observations, focusing on estimation of restricted mean survival times using a quadratic loss function. A drawback of using pseudo observations in prediction is that it can give risk estimates that are negative or greater than 1. The prediction model also has to be estimated separately for different time horizons $\tau$, which can lead to survival probabilities that not monotonically decreasing over time.
@@
In our implementations of the methods our aim was not to conduct a formal comparison, but instead to illustrate how they can be applied. Some simulation based comparisons of the methods discussed in this paper have been conducted elsewhere (Munch and Gerds, 2025; Wolock and others, 2024). We considered using both a relatively simple set of candidate learners and a richer set of learners. In practice, using a diverse set of learners is a good idea, as it results in lower correlation between the errors from different models, which is desirable (Brown and others, 2005). For example, Wey and others (2015) suggested that including lots of similar models (e.g different Cox models, with different ways of entering the covariates) is not advisable, as the predictions from these models will be highly correlated. An overview of machine learning algorithms for survival data is provided by Sonabend and Bender (2025). In practice, the choice between different SL methods, and the choice of candidate learners to be included, is likely to be based on practical considerations, including computational.
We have placed an emphasis on the use of the SL as a tool for developing a prediction model. However, as noted earlier, estimation of conditional survival and censoring distributions is also a component of estimation procedures used in causal inference. For example, suppose that the causal estimand of interest is a marginal survival probability under a particular treatment strategy. In targeted maximum likelihood estimation (TMLE) the conditional survival and censoring distributions are nuisance models that need to estimated as part of a series of analysis steps used to estimate such a causal estimand, and the SL is typically used to estimate the nuisance models (Moore and van der Laan, 2011). Westling and others (2023) also used their SL approach to estimate nuisance models used as part of another type of doubly-robust estimation procedure for marginal survival curves under a particular treatment strategy. The methods described in this tutorial are therefore also relevant for those wishing to gain insights into the SL as used as a tool in causal inference estimation methods.
@@
Table S1: Coefficients (or weights) for the candidate learners (based on a simple set of candidate learners) for discrete-time SL, for 10 or 100 time periods and using the $L2$ or loglikloglik loss function.
Candidate learner 10 periods 10 periods 100 periods 100 periods
$L2$ loss loglikloglik loss $L2$ loss loglikloglik loss
@@
Table S3: Coefficients (or weights) for the candidate learners (based on a richer set of candidate learners) for discrete-time SL, using 10 time periods and using the $L2$ loss function.
@@
Discrete-time SL: 100 periods, $L2$ loss 0.213 13.5 72.1 75.3
Discrete-time SL: 100 periods, loglikloglik loss 0.201 18.3 71.9 74.8
Discrete-time SL: 100 periods, IPCW loss 0.224 9.1 70.9 73.7
S1 SL based on pseudo observations